(a) Assume that although Marrow firm is presently making accounting profit it is experiencing negative economic profit in its current venture. The owner of the firm does not understand what this means and asks


you to explain to him.


(b) Massy super store, wishes to increase its total revenue. It offers 30% discount on every item it sells.


Group A


1.79 million


1.84 million


Group B


1.76 million


1.86 million


Miley Cyrus - The C.


cofi


G


G sermon about pra...


Never Knew Am Pr.


of


estion


Volume of sales before the


30% discount


Volume of sales after the


30% discount


(a) Using the midpoint method, calculate the price elasticities of demand for group A and group B. 4 Marks


(b) If Massy wants to increase its total revenue, which group should it offer the discount to?


4 Marks


1
Expert's answer
2021-03-08T07:18:27-0500

(a) Because when we calculating economic profit we subtracts from the total revenue explicit costs and implicit costs. The last one includes the opportunity costs – the potential benefits foregone when an option is not chosen. Since the opportunity costs do not include in the accounting profit Marrow firm is presently making accounting profit. But the economic profit is negative. The economic profit used to best assist management with decision-making. Therefore, the owner and the management of the firm can analyze the previous decisions and increase the accounting profit of the firm in the future.

(b) The price elasticity of demand using the midpoint method can be calculated as follows:


"E_d=\dfrac{\%\Delta Q}{\%\Delta P}=\dfrac{\dfrac{Q_2-Q_1}{(Q_2+Q_1)0.5}}{\%\Delta P},""E_{dA}=\dfrac{\dfrac{1.84-1.79}{(1.84+1.79)\cdot0.5}}{-0.3}=-0.092,""E_{dB}=\dfrac{\dfrac{1.86-1.76}{(1.86+1.76)\cdot0.5}}{-0.3}=-0.184."

(c) As we can see from the calculations, in both cases the demand is inelastic (the price elasticity of demand is less then 1). In the case of inelastic demand, the demand is unresponsive to the change in price. In other words, if the price for an inelastic good is lowered by 30%, the demand for that good doesn't increase, resulting in less overall total revenue due to the lower price and no change in demand. Therefore, if the Massy wants to increase its total revenue, it should don't offer the discount to any of the group (A or B), because after that its total revenue will decrease.